Grade 12Physical Chemistry

What is the iupac name of [co (nh3)5 (co3)]cl with oxidation number?

elan
8 Years agoGrade 12

4 Answers

Arun
8 Years ago
Dear student
 
Pentaaminecarbonatocobalt(III)chloride
 
For oxidation no of Co
x + (5*0) +(-2) +(-1) = 0
x = +3
 
Regards
Arun (askIITians forum expert)
Sammy
7 Years ago
It is pentaamminecarbonatocobalt(III)bromide you used only single m instead of mm please do not do like this
Kushagra Madhukar
5 Years ago
Dear student,
Please find the solution to your problem below.
 
Given, [Co(NH3)5(CO3)]Cl
For oxidation no of Co
x + (5*0) +(-2) +(-1) = 0
x = +3
Hence the name of the compound is
Pentaamminecarbonatocobalt(III)Chloride
 
Hope it helps.
Thanks and regards,
Kushagra
Abhishek Sharma
4 Years ago
X+0+(-2)+(-1) =0
X-2-1 =0
X=+3
Okk I submitte my answer you calculate it again and mention mediate colculation
You practice so many questions like this.
Then you able to colculate oxidation state of that's type of questions.